OpenStudy (anonymous):

2f/7=f-2/14 the slashes are fraction bars

OpenStudy (anonymous):

multiply both by 14. we get 4f = f-2 3f = -2 f=-2/3

OpenStudy (anonymous):

2f/7-f=-2/14 -5f/7=-2/14 f=1/5

OpenStudy (anonymous):

which one is right

OpenStudy (anonymous):

ahh.. put brackets to make it more clear please lol

OpenStudy (anonymous):

he has not multiplied f with14

OpenStudy (anonymous):

oh thanks dipankarstudy can you help me with my other question

OpenStudy (anonymous):

according to gg... 4f=14f-2 -10f=-2 f=1/5.... same as mine...

OpenStudy (anonymous):

thanks ggabnore anyway

OpenStudy (anonymous):

put your questions....i will try

OpenStudy (anonymous):

Cost of dinner for a party of eight people is 139 so for large groups of people an 18% gratuity is added to cost of dinner after a 6% sales tax. Find the total bill for the dinner.

OpenStudy (anonymous):

i need to make sure gabners answers right

OpenStudy (anonymous):

how? no i read the queestion as 2f/7 = (f-2)/14 thats why i got my answer lol

OpenStudy (anonymous):

Cost of dinner for a party of eight people is 139 so for large groups of people an 18% gratuity is added to cost of dinner after a 6% sales tax. Find the total bill for the dinner.

OpenStudy (anonymous):

yes i think he is right thts why i have not answered.....

OpenStudy (anonymous):

do u have other ans...?

OpenStudy (anonymous):

oh ok good job ggabnore

OpenStudy (anonymous):

ya one more question

OpenStudy (anonymous):

of course....

OpenStudy (anonymous):

ill tyoe it as a question right now

OpenStudy (anonymous):

come to my other question